If you pay a membership fee in a concert ticket purchasing club, you can buy 5 tickets for $242 (total cost of membership and 5 tickets) or 10 tickets for $427 (total cost if membership and10 tickets). How much is the membership fee?
Answer by stanbon(75887)

If you pay a membership fee in a concert ticket purchasing club, you can buy 5 tickets for $242 (total cost of membership and 5 tickets) or 10 tickets for $427 (total cost if membership and10 tickets). How much is the membership fee?
------
Equations:
m + 5t = 242
m +10t = 427
----
Subtract and solve for "t":
5t = 185
t = $37 (pice of each ticket)
----
Solve for "m":
m + 5t = 242
m + 5*37 = 242
m = 242-185
m = $57 (membership fee)
=========================
Cheers,
Stan H.
===================